### v1

# Clustering is back: Reaching state-of-the-art LiDAR instance segmentation without training

Abstract
Panoptic segmentation of LiDAR point clouds is fundamental to outdoor scene understanding, with autonomous driving being a primary application. While state-of-the-art approaches typically rely on end-to-end deep learning architectures and extensive manual annotations of instances, the significant cost and time investment required for labeling large-scale point cloud datasets remains a major bottleneck in this field. In this work, we demonstrate that competitive panoptic segmentation can be achieved using only semantic labels, with instances predicted without any training or annotations. Our method achieves performance comparable to current state-of-the-art supervised methods on standard benchmarks including SemanticKITTI and nuScenes, and outperforms every publicly available method on SemanticKITTI as a drop-in instance head replacement, while running in real-time on a single-threaded CPU and requiring no instance labels. Our method is fully explainable, and requires no learning or parameter tuning. Code is available at https://github.com/valeoai/Alpine/
1 Introduction
Semantic
predictions
Instances
predictions
BEV projection
Radius-based
graph construction
Connected
component extraction
Figure 1: ALPINE clustering. For a given semantic class a) we project the points in the BEV space (subsampled on the figure for visualization purpose), b) we build a kNN graph and filter by edge length and c) we extract the connected components.
To move autonomously in an outdoor environment, an agent must understand and segment its surroundings into categories. With a LiDAR, it involves recognizing the semantics of points and identifying individual instances of “things” (e.g., ’cars’ or ’pedestrians’). This capability, known as LiDAR panoptic segmentation, is crucial for tasks such as object avoidance, trajectory forecasting, and path planning.
State-of-the-art approaches typically combine semantic segmentation and instance segmentation by training end-to-end networks using query-based mechanisms [35, 70, 76, 19] or regressing instance centers [30, 31, 37]. Their architectures usually involve instance prediction heads, which require panoptic labels for training.
Instead, early methods  [24, 44, 7, 79, 84] proposed unsupervised, learning-free, clustering algorithm to extract instances, sometimes completed by semantic segmentation networks to obtain panoptic segmentation [84]. These methods appear to perform poorly in panoptic segmentation benchmark. One is easily tempted to disregard them and prefer to use end-to-end panoptic segmentation networks requiring costly panoptic segmentation labels.
This paper demonstrates that achieving state-of-the-art panoptic segmentation does not require end-to-end networks or panoptic labels. The early pipeline combining semantic segmentation and unsupervised clustering suffices, as validated by extensive comparisons with current LiDAR panoptic methods. We achieve this result thanks to the use of modern, high-performing semantic segmentation networks and to a clustering algorithm inspired by [33] with key changes to boost its performance.
This work has several practical consequences: (a) There is no need for instance labels, which are in any case cumbersome to create; only class labels for training semantic segmentation are required. (b) The architecture is simpler since it does not require a trainable instance head. (c) The method can be directly applied on top of any semantic prediction, with no adaptation or required access to underlying features. (d) As the clustering algorithm is extremely fast, the panoptic segmentation can be produced at high frequency, as required for embedding on an autonomous vehicle, while using very little compute resources.
Refer to caption Refer to caption Refer to caption
D&M [84] w. MinkUNet ALPINE w. MinkUNet Ground Truth
Figure 2: Examples of Instance predictions on SemanticKITTI. We present the results obtained with D&M [84] and ALPINE restricted to the car class. Both methods are training-free clusterings, and use the same MinkUNet to obtain pointwise semantic predictions. When compared to the Ground Truth, we notice that D&M does not satisfactorily separate the cars while ALPINE segment them correctly.
This clustering, which we call ALPINE for “A Light Panoptic INstance Extractor”, works by constructing a nearest-neighbor graph on semantically consistent points, followed by some refinement using a box splitting mechanism. ALPINE builds on top of any off-the-shelf semantic segmentation network. It is a learning-free method that does not require any knowledge of instance labels relying instead on a rough estimate of per-class object size, which is set using publicly available information found on the web. That makes ALPINE a tuning free method, easy to use on any LiDAR dataset given a semantic segmentation backbone.
2 Related Work
Figure 3: Overview. In ALPINE we take the output of a semantic segmentation model and apply our clustering algorithm on each things classes to obtain instance masks and form panoptic predictions.
2.1 Semantic segmentation
Semantic segmentation permits us to separate, at the point level, the main types of objects in a scene. Several methods and backbones have been designed for LiDAR semantic segmentation. They are usually classified under four main categories: point-based [46, 47, 48, 13, 45] with the most recent ones leveraging a transformer architecture [82, 68, 67], projection-based [38, 3, 80, 4, 25, 71], sparse convolution-based [14, 87, 29, 12], fusion-based when leveraging different point cloud representations [17, 61, 72, 49].
We present a method that requires semantic predictions as input, regardless of how they are obtained. We test our method using different architectures for semantic segmentation: MinkUNet [14], WaffleIron [45], PTv3 [68]. We chose these methods for their competitive performance and diverse architectures: voxel-based using sparse convolutions, point-based using successive projections on 2D planes, point-based using attention layers.
2.2 Panoptic segmentation
Panoptic segmentation is an extension of semantic segmentation where one should also segments all instances of objects in “thing” classes.
Detection-based methods.
The first methods for panoptic segmentation were built on top of networks for semantic segmentation and object detection. The object detection branch puts boxes around each object instance, and the semantic segmentation branch is used to obtain the semantic class of each point. Baseline methods are constructed using off-the-shelf detector [26, 57, 75] and segmentor [64, 61, 14, 12]. More efficient methods often train the two tasks jointly  [58, 22, 74, 77, 2], with additional losses specifically designed for panoptic segmentation.
Clustering-based methods.
These type of methods also appeared quite early. Our method actually falls in the category. These methods works in two successive steps [83]. A semantic segmentation backbone is first used to assign a class to each point. The points labeled as “thing” are then clustered to find instances.
One of the earliest method, the Euclidean Cluster [83, 54, 53] consists in sequentially visiting each point, assigning to this point a new instance index if it is not already labeled, and labeling all its unlabeled neighbors within a certain radius with the same instance index. Another proposition consists in adapting SLIC [1] from images to point clouds, using voxel representations [44, 83]. More advanced methods [79, 84] design a clustering method tailored for LiDAR point cloud, by taking into consideration the scan-line nature of the sensor’s capture. The hyperparameters of these methods are typically tuned according to expected object sizes and/or sensor properties.
More recent methods  [20, 21, 86, 30, 39, 27, 16, 31, 36, 81, 69, 32, 73, 81] leverage panoptic annotations and replace some hand-crafted steps of the clustering process by learnable steps. For example, GP-S3Net [51] creates a graph for points of thing classes using learned point features to obtain edge features, then perform edge classification.
Among clustering-based methods, a common approach consists in regressing for each point the position of the corresponding instance centroid, allowing an easier and less ambiguous clustering done on the predicted cluster centers [20, 21, 28, 37, 31], while having an easy to define training objective. This approach is refined in Panoptic-PHNet where centroid-offset prediction and clustering are performed in the Bird’s-Eye-View (BEV) to mitigate vertical over-segmentation.
Query-based methods.
These methods [18, 35, 76, 70, 19] build upon MaskFormer [11], an image segmentation method. To propose an end-to-end unified framework for panoptic segmentation, they create a set of (usually learnable) queries fed into a transformer decoder in charge of (softly) assigning points and queries. Each query encodes a different instance. Points responding maximally to the same query are part of the same instance. In these architectures, the point features given to the transformer decoder can be obtained from different backbones working on range-view [18], voxels [35], BEV [76], or using both cartesian and polar space [70] representation of the point cloud.
Use of annotations.
Some datasets offer annotations for multiple tasks such as semantic & panoptic segmentation and object detection [8]. Some methods leverage all the available annotation and train a single model on all tasks jointly, benefiting from the additional supervision [77, 10]. At the other end of the spectrum, some methods [41, 42, 56] segment objects without using any annotation. However, these method are unable to predict semantics. In this work, we take inspiration from early clustering-based methods using only semantics annotations.
3 Method
Our approach is designed around three main components: a fast clustering algorithm having its roots in RBNN [24] and LESS [33], an annotation-free and training-free parameter selection process, and a refinement mechanism based on box splitting.
3.1 ALPINE clustering
The overview of our instance clustering process is depicted in Fig. 1. It is inspired by the clustering done in [33] and works in four main steps.
One semantic class at a time.
After semantic prediction, we consider each “thing” class $c$ independently. We denote by $P_{c}$ the original point cloud $P$ restricted to the points with predicted label $c$. We denote $n_{c}$ the size of $P_{c}$. Each point cloud $P_{c}$ is processed separately.
Projection in BEV.
Inspired by previous works [26, 55, 30] leveraging the fact that objects of “thing” classes are rarely stacked vertically, we project the 3D points in the Bird’s Eye View (BEV) representation (orthogonal projection onto the $(x,y)$ plane). In the following, we assume that points clouds $P_{c}$ are therefore two-dimensional.
$k$-nearest neighbors graph.
For each point in $P_{c}$, we obtain the list of its $k$ closest neighbors (typically $k=32$), using a 2D-Tree [6]. The distances between the points are computed using only the point coordinates in the BEV representation, and therefore do not involve any learned features. We then construct a directed graph $\mathcal{G}$ where the nodes are the points in $P_{c}$ and each point is connected by an edge to its $k$ nearest neighbors, at most. We remove connections between points if the distance in BEV is larger than a threshold (see Sec. 3.2). Finally, the adjacency matrix of this directed graph is made symmetric by adding missing half-edges to obtain an undirected graph. The choice of $k$ is not very critical for the method; it needs to be large enough not to miss too many connections with isolated points, while keeping it small enough to benefit from fast runtime. This is verified experimentally in Sec. 4.6. The complexity of the construction of the 2D-Tree is $\mathcal{O}(n_{c}\log n_{c})$, while the search for the $k$ nearest neighbors for all points is $\mathcal{O}(n_{c}\,k\log n_{c})$.
Connected components.
The clusters themselves are obtained by identifying the connected components in the constructed graph. The complexity of this algorithm is $\mathcal{O}(n_{c}k)$ in the worst case. We associate to each component a different instance index.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Example of bounding box splittings. In the top examples, two cars are parked close to each other. While merged by the clustering, they do not fit in a car’s box and the cluster is then split. In the bottom examples, the same mechanism is applied to two pedestrians in a bus shelter. Boxes are shown in 3D for illustrative purposes but the mechanism is purely 2-dimensional.
Key differences with LESS and RBNN
ALPINE builds upon the strong graph-based clustering foundations laid by RBNN [24] and LESS [33]; however applying a few key changes making it both stronger and more practical to use. Unlike in both other methods, the adjacency of each point is limited to at most $k$ reducing the complexity of the algorithm from $\mathcal{O}(n_{c}^{2})$ to $\mathcal{O}(n_{c}\,k\log n_{c})$, scaling better to large clusters; it is also applied class-wise, in BEV, with a per-class constant threshold. Besides, while conceptually similar, RBNN did not use connected components to find clusters, relying instead on a custom cluster merging scheme.
3.2 Choice of distances and thresholds
In order to decide whether two points should be connected in $\mathcal{G}$, and thus should belong to the same instance, we use a threshold $t_{c}$ on the distance between two points in BEV.
Class-wise thresholds.
When considering whether two points should be connected, an important quantity to consider is the typical size of the objects to segment: points that are $3$ meters apart are never going to belong to the same pedestrian, but might belong to the same truck. We therefore use separate thresholds $t_{c}$ for each semantic class $c$. The only information that we require to set $t_{c}$ is the an estimate of the average size of the objects in each class.
This information can be obtained without any ground-truth annotation, and even without any access to the dataset. For example, for cars, one can retrieve from the Internet the average size of vehicles where the system is to be deployed. For each class, we can thus get an estimate of the typical size of bounding boxes around objects in this class. We set $t_{c}$ to the smallest side of this reference bounding box. More considerations as to how we obtained such approximate average boxes are given in Sec. 4.2.
This parametrization addresses a known domain shift issue related to sizes when training and testing in different countries [65]. It enables a simple form of domain adaptation where only a few scalars $t_{c}$ have to be set from one country to another, besides any adaptation or generalization regarding the underlying semantics.
Range vs threshold.
The distance between two points in a LiDAR point clouds naturally increases with their range (distance between a point and the sensor): LiDAR point clouds are less dense far away from the sensor. For this reason, to similarly create object clusters and decide if an edge between two points should be removed or not, LESS [33] used a threshold that varies with the range. We tested such a strategy in Sec. 4.6 but found no improvement in doing so. We thus simply use a constant class-wise threshold.
3.3 Box splitting
To refine our segmentation further, we use a box-splitting strategy. When a cluster of a given semantic class $c$ does not fit within the reference bounding box of this class (as defined in Sec. 3.2), enlarged by a constant (not class-wise) proportional margin (of 30%) to account for bigger-than-average objects, it likely actually contains multiple instances of this class. In this case, our box-splitting method performs a binary search on the threshold parameter $t$ to recursively find the largest clusters that all fit into the reference bounding box. The pseudo-code for this algorithm is given in Alg. 1, and an illustrative example is provided in Fig. 4. The box-fitting procedure that is used internally in the algorithm is borrowed from the literature [78] and discussed in the appendix.
1 function split_cluster(P, B, t)
    input: Points P, Average box B, Threshold t
    output: List of cluster points
2    if box_fitting(P)fits in B then
        return [P] // Points already fit in B
3
4   else
5        $\texttt{t}\leftarrow\texttt{t}/2$
[…]
7        while true do
8            $\texttt{dt}\leftarrow\texttt{dt}/2$
[…]
10            if len(C) = 1  then
[…]
11
12           else if len(C) > 2  then
[…]
13
14           else
15                return split_cluster(C[0], B, t) + split_cluster(C[1], B, t)
16
17
18
Algorithm 1 Box splitting algorithm. This algorithm splits clusters that do not fit in $B$ into sub-clusters that all fit in $B$. box_fitting returns the smallest bounding box encompassing $P$ and clusterize is the clustering algorithm described in Sec. 3.1.
4 Experiments
{NiceTabular}
[…]
lbs.
LPSAD (in [51]) [39] 37.4 44.2 47.8 66.9 49.4
Panop.TrackNet [22] 40.0 - 48.3 73.0 53.8
PointGroup [23] 46.1 54.0 56.6 74.6 55.7
TORNADO-Net [17] 50.6 55.9 62.1 74.9 59.2
Panoster [16] 55.6 - 66.8 79.9 61.1
LCPS (lidar) [81] 55.7 65.2 65.8 74.0 61.1
D&M (Cyl.3D) [84] 57.2 - - - -
Location-Guided [69] 59.0 63.1 69.4 78.7 61.4
Panop.-PolarNet [86] 59.1 64.1 70.2 78.3 64.5
EfficientLPS [58] 59.2 65.1 69.8 75.0 64.9
MaskPLS [35] 59.8 - 69.0 76.3 61.9
DS-Net (SPVCNN) [21] 61.4 65.2 72.7 79.0 69.6
Panop.-PHNet [30] 61.7 - - - 65.7
PANet [36] 61.7 66.6 71.8 79.3 68.1
CenterLPS [37] 62.1 67.0 72.0 80.7 68.1
P3Former [70] 62.6 66.2 72.4 76.2 -
CFNet [31] 62.7 67.5 - - 67.4
LPST [2] 63.1 70.8 73.1 79.2 69.7
GP-S3Net [51] 63.3 71.5 75.9 81.4 73.0
DQFormer [76] 63.5 67.2 73.1 81.7 -
ALPINE (Mink.) 64.2 68.9 74.1 84.4 70.7
PUPS (w/o ens.) [60] 64.4 68.6 74.1 81.5 -
IEQLPS [19] 64.7 68.1 74.7 81.3 -
Methods with Test-Time Aug. and/or model ensembling
D&M (Mink.) [84] (✓) 61.8 66.2 72.8 79.6 71.4
ALPINE (PTv3) (✓) 62.4 66.2 72.0 76.7 67.3
ALPINE (WI) ** (✓) 64.2 69.0 74.1 79.7 70.3
ALPINE (Mink.) (✓) 65.9 70.2 75.5 81.4 72.2
PUPS [60] 66.3 70.2 75.6 82.5 -
ALPINE (Mink.+PTv3)   (✓) (✓) 66.6 70.8 76.1 82.6 72.0
Mink.: MinkUNet [14]        WI: WaffleIron-256 [45]
*: the publicly available model for WI-256 [45] used instance
    annotations in its data augmentation pipeline
Table 1: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only. The main metric is the PQ.
{NiceTabular}
@l@l—@ c@ c@ c@ —cccc—c@
[…]
lbs.
LPSAD (in [51]) [39] 50.4 57.7 62.4 79.4 62.5
Panop.TrackNet [22] 51.4 56.2 63.3 80.2 58.0
VIN [85] 51.7 57.4 61.8 82.6 73.7
TORNADO-Net [17] 54.0 59.8 65.4 80.9 68.0
MaskPLS [35] 57.7 60.2 66.0 71.8 62.5
GP-S3Net [51] 61.0 67.5 72.0 84.1 75.8
EfficientLPS [58] 62.0 65.6 73.9 83.4 65.6
DS-Net(SPVCNN) [21] 64.7 67.6 76.1 83.5 76.3
PVCL [32] 64.9 67.8 77.9 81.6 73.9
SCAN [73] 65.1 68.9 75.3 85.7 77.4
Panop.-PolarNet [86] 67.7 71.0 78.1 86.0 69.3
SMAC-Seg [27] 68.4 73.4 79.7 85.2 71.2
PANet [36] 69.2 72.9 80.7 85.0 72.6
CPSeg [28] 71.1 75.6 82.5 85.5 73.2
LCPS (lidar) [81] 72.9 77.6 82.0 88.4 75.1
PUPS [60] 74.7 77.3 83.3 89.4 -
Panop.-PHNet [30] 74.7 77.7 84.2 88.2 79.7
CFNet [31] 75.1 78.0 84.6 88.8 79.3
P3Former [70] 75.9 78.9 84.7 89.7 -
CenterLPS [37] 76.4 79.2 86.2 88.0 77.1
ALPINE (WI) 76.9 79.9 85.7 89.3 80.3
IEQLPS [19] 77.1 79.1 86.5 88.2 -
LPST [2] 77.1 79.9 86.5 88.6 80.3
DQFormer [76] 77.7 79.5 86.8 89.2 -
Methods with Test-Time Aug. and/or model ensembling
ALPINE (WI) (✓) 77.9 80.7 86.5 89.6 81.4
ALPINE (PTv3) (✓) 78.9 81.3 87.0 90.4 81.5
ALPINE (WI+PTv3)   (✓) (✓) 79.5 81.9 87.6 90.5 82.7
LidarMultiNet [77] 81.8 - 89.7 90.8 83.6
WI: WaffleIron-768 [45]
Table 2: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only. The main metric is the PQ.
{NiceTabular}
@l@l—@ c@ c@ c@ —cccc—c@
[…]
lbs.
LPSAD (in [51]) [39] 22.5 32.7 34.0 53.5 35.5
TORNADO-Net [17] 33.7 43.3 46.0 68.4 44.5
DS-Net [20] 35.6 45.9 49.2 68.6 54.5
GP-S3Net [51] 48.7 60.3 63.7 61.3 61.8
ALPINE (PTv3) 51.4 57.7 67.2 74.3 58.3
Table 3: Panoptic segmentation results on Sequence $02$ of SemanticPOSS as validation set. The main metric is the PQ.
4.1 Datasets and metrics
nuScenes [8] is a dataset captured in Boston and Singapore, with a $32$-beam LiDAR sensor. It contains $10$ things classes and $6$ stuff classes. The instance labels were automatically generated by combining object detection and semantic segmentation labels, and were not manually curated. The results are reported on the official validation set.
SemanticKITTI [5] is a dataset captured around Karlsruhe in Germany with a 64-beam LiDAR sensor. It contains $8$ things classes and $11$ stuff classes. The instance labels were manually annotated. The results are reported on Sequence 08, commonly used as the validation sequence.
SemanticPOSS [43] is a smaller dataset captured in the surroundings of the Peking University with a $40$-beam LiDAR, focusing on having many dynamic (thing) objects. It contains only $3$ things classes, namely person, rider and car, as well as $10$ stuff classes. The results are reported on Sequence 02, used as a validation sequence as done in [51].
We use standard metrics, which are defined in Appendix.
4.2 Size parameters setting
The parameters needed in ALPINE, as described in Secs. 3.2 and 3.3, are set after estimating the average size of object bounding boxes in each class.
We consider two kinds of size parameter settings: the web-based setting obtains sizes from a few queries on the Internet; the dataset-based setting exploits ground-truth knowledge coming with datasets.
For web-based box sizes on nuScenes, we used the publicly available average size of cars in the United States in 2018 (15.6 $\times$ 6.3 ft) [15]; we made the simplifying assumption that buses, trailers, trucks and construction vehicles have size 10 $\times$ 3 m; we used standard bike and motorcycle sizes found online [63, 9] ; we made the the approximation that pedestrians fit in box of half their arm span, estimates using [50, 66]; and we made an educated size guess of 2 $\times$ 0.5 m for barriers and 40 cm for cones.
Similar web-based information was exploited to set the sizes for SemanticKITTI. We used the average car size in Europe of 4.4 $\times$ 1.8 m [40]. The same found sizes for bicycle and motorcycle as in nuScenes [63, 9] as well as for bicyclist and motorcyclist. The class person was treated as in nuScenes. For ’truck’ and ’other-vehicle’ we made the same assumption as in nuScenes of a 10 $\times$ 3 m size.
For the $3$ classes of SemanticPOSS, we used the value we obtained for the equivalent class in SemanticKITTI, i.e. pedestrians, bicyclists and cars.
For dataset-based box sizes, we use the average size of the bounding boxes used in [62, 75]. On nuScenes, these average size is provided for all $10$ thing classes. On SemanticKITTI, these sizes are only available for cars, pedestrians and bicyclists; we use the value computed on nuScenes for the 5 remaining classes. On SemanticPOSS, we just re-use the sizes obtained on SemanticKITTI.
In the end, we did not witness a significant change of performance between both sets of parameters in the experiments that follow and use by default, in all results, the annotation-free web-based sizes.
4.3 Comparison to the state-of-the-art
sem. Original head ALPINE FPS
[…]
DS-Net 63.5 57.7 63.4 61.8 58.5 64.1 63.6 0.7
Panoptic-P.Net 64.4 58.9 63.9 65.2 59.5 64.5 66.7 1.6
MaskPLS 61.9 59.8 63.6 63.7 60.2 64.0 64.7 N/A
LCPS (lidar) 64.3 61.3 65.6 70.0 61.4 65.6 70.1 1.8
PANet 67.9 61.2 66.1 66.7 61.3 66.3 66.9 7.1
P3Former 66.8 62.6 66.2 69.4 62.9 66.5 70.1 N/A
Table 4: Comparison between ALPINE and the instance segmentation capabilities of other methods, on the validation set of SemanticKITTI. For each panoptic segmentation method with available code and checkpoint, we isolated the plain semantic predictions and report the mIoU (‘sem.’), and constructed on top of them our instances (‘ALPINE’), which we compare to the instances obtained by the original methods (‘Original head’). An indicative instance prediction frequency (FPS) is given after subtracting the time of the semantic forward pass to the total runtime.
On all datasets, we apply ALPINE combined with different backbones trained for semantic segmentation. We re-used the provided checkpoints when publicly available or retrain them ourselves with the provided code.
SemanticKITTI.
Sec. 4 presents our results on the validation set of SemanticKITTI. On this dataset, we consider 3 different semantic segmentation backbones: PTv3 [68], WI-256 [45], and MinkUnet [14] from OpenPCSeg [34].
ALPINE is able to reach SOTA performance (PQ), outperforming, e.g., the end-to-end method PUPS [60]. Moreover, when comparing to other methods at equivalent semantic segmentation mIoU, ALPINE is comparable or outperforms these methods: ALPINE w. PTv3 vs CFNet [31], PANet [36], CenterLPS [37]; or ALPINE w. MinkUNet vs GP-S3Net [51]. It shows that ALPINE benefits easily from better semantic segmentation backbones.
nuScenes.
Sec. 4 presents our results on the validation set of nuScenes. On this dataset, we consider two different semantic segmentation backbones: PTv3 [68] and WI-768 [45]. We are only outperformed by LidarMultiNet [77], which was trained using both semantics and object boxes, using ensembling and test-time augmentations (TTA) on both tasks, and benefits from a better semantic prediction.
SemanticPOSS.
Sec. 4 presents results obtained on Sequence 02 of SemanticPOSS. We notice that ALPINE combined with PTv3 [68] reaches state-of-the-art results.
ALPINE $\dagger$ DBSCAN HDBSCAN D&M
PQ FPS PQ FPS PQ FPS PQ FPS
65.5 14.4 Hz 56.7 3.2 Hz 55.1 4.5 61.8 0.5 Hz
[…]
4.4 Study of ALPINE’s components
Comparison to other instance heads.
Tab. 4 shows how all panoptic methods for which we found code and checkpoints on SemanticKITTI compare against ALPINE at equal semantic prediction: we use the semantic prediction of the original method but replace the instance prediction mechanism with ALPINE. We show that ALPINE outperforms every single method, while using no annotation and requiring no GPU. This further demonstrates that instance prediction does not need complex design, training or instance annotations to be competitive. An indicative “instance runtime” of all methods have been obtained by subtracting the segmentation forward pass to the total runtime, when this was possible. This approximate measurement is comparable (while not directly equivalent) to the ALPINE FPS reported in Tab. 5.
Comparison to other clustering methods.
Tab. 5 shows that ALPINE is both faster and better performing on our task than other clustering methods, namely the standard DBSCAN and HDBSCAN, as well as D&M [84], which was specifically designed for this task too. We are also within the limit of real-time processing, as SemanticKITTI is captured at 10 Hz, while we use only limited computing resources (single-core CPU usage). In these experiments, all clustering methods use the same semantics, predicted by MinkUNet [14]. The box splitting was not used on any method in order to compare clustering capacity only. DBSCAN and HDBSCAN were performed on BEV, with fixed parameters of $\epsilon=1$ for DBSCAN and parameters gathered from 3DUIS [41] and $\epsilon$ proportional to $t_{c}$ for HDBSCAN; with the proportionality coefficient obtained with a fine parameter search. Those settings were those that we found worked best. We see that our clustering is outperforming D&M by a comfortable margin: $+3.7$ PQ pts. With box-splitting, ALPINE would gain $+0.4$ further PQ pts.
Method PQ RQ SQ mIoU
3DUIS 60.2 70.7 79.7 72.2
UNIT 61.1 72.0 79.0 72.2
ALPINE 65.9 75.5 81.4 72.2
Table 6: Comparison to semantic-free instance segmentations evaluated on the SemanticKITTI validation set.
Comparison to semantic-free instance segmentation
3DUIS [41] and UNIT [56] are two unsupervised instance segmentation methods. They propose to find individual instances of objects in a totally unsupervised manner, regardless of semantics. While their task is more involved, as they cannot rely on predicted semantics mask, we can still compare their instance predictions with that of ALPINE. In practice, we use the intersection of their instance mask and each semantic mask as panoptic predictions. As can be noted in Tab. 6, ALPINE significantly outperforms those methods. It can still be noted that both unsupervised method work quite impressively, almost reaching the level of performance of D&M while not using any semantic information.
Sem. Seg Inst. Seg PQ RQ SQ mIoU
Oracle ALPINE 96.2 98.4 97.8 100.0
WI-768 Oracle 81.7 88.1 94.1 81.4
PTv3 Oracle 83.2 87.6 92.8 81.5
Oracle Oracle 100.0 100.0 100.0 100.0
Table 7: Oracle results evaluated on the nuScenes validation set.
4.5 Evaluation of the clustering upper bound
To evaluate what remains to be gained by developing better semantics or instance heads, we devise an oracle head for both subtasks: a “semantic oracle” using the ground-truth semantic segmentation as semantic predictions, and an “instance oracle” splitting the semantic predictions of a given method using the ground-truth instance boundaries, without affecting the semantic prediction. Results, in Tab. 7, shows how much scores are expected to improve with both oracles. When using the “semantic oracle”, the PQ on nuScenes raises to 96.2, (and 99.0 on SemanticKITTI), while using PTv3 predictions and the instance oracle only raises the PQ by +4.3 to 83.2, showing that instance extraction is much more saturated than semantic segmentation.
[…]
splitting
const. 76.3 78.7 81.5
const. 76.9 79.3 81.5
const. 77.1 79.5 81.5
const. 78.9 81.3 81.5
$\propto|d|$ [33] 75.9 78.3 81.5
Table 8: Ablation study of our different clustering components on the nuScenes validation set. $|d|$ is the distance to the sensor.
Method sem. statistics-based web-based
[…]
SemanticKITTI
MinkUNet w/o TTA 70.7 64.1 68.8 72.5 64.2 68.9 72.6
PTv3 67.3 62.4 66.1 65.9 62.4 66.2 66.1
WI-256 70.3 64.0 68.9 69.0 64.2 69.0 69.3
MinkUNet 72.2 65.8 70.1 73.7 65.8 70.1 73.7
MinkUNet & PTv3 72.0 66.5 70.7 73.2 66.6 70.8 73.4
nuScenes
WI-768 w/o TTA 80.3 77.2 80.1 78.3 76.9 89.9 77.9
WI-768 81.4 78.2 81.0 79.0 77.9 80.7 78.6
PTv3 81.5 79.0 81.4 79.6 78.9 81.3 79.3
WI-768 & PTv3 82.7 79.8 82.2 80.5 79.5 81.9 80.2
SemanticPOSS
PTv3 58.3 51.1 57.4 63.6 51.4 57.7 64.8
Table 9: Comparison between statistics-based and web-based parameters settings. More run details can be found in Secs. 4 and 4.
4.6 Ablation and sensitivity study
Influence of each ALPINE modules.
We present our ablation study in Tab. 8, conducted on the validation set of nuScenes, using PTv3 as semantic backbone. We see that constructing the graph in BEV or post-processing the result with our box splitting strategy improve the results: respectively $+0.6$ point and $+0.8$ point in PQ over the baseline. The combination of both improve the PQ by $2.6$ points.
Instead of using a constant threshold to keep or remove the edge between two points when constructing our graph, we also tested a threshold varying linearly with the range, as used in LESS [33]. The linear coefficient is the ALPINE threshold $t_{c}$ multiplied by a constant optimized at dataset scale to maximize performance. Despite this optimization, we find that, in our case, using a constant threshold depending on the typical size of the objects leads to better results.
Dataset-based vs. web-based results.
Tab. 9 gives the difference in scores for the different sets of parameters described in Sec. 4.2. Both choices of parameters give very similar results, from which we conclude ALPINE parameters are both easy to setup, have a sensible meaning, and do not require annotations.
16326412860606060656565657070707065.665.665.665.665.965.965.965.966.066.066.066.066.066.066.066.0PQk
Figure 5: Influence of the neighbors count $k$ evaluated on SemanticKITTI’s validation set
Number of neighbors $k$.
We study the variation of the performance of ALPINE with the choice of the number of neighbors considered in the clustering methods $k$. The results are presented in Fig. 5. As mentioned in Sec. 3.1, the choice of $k$ is not critical above a certain value as edges to distant neighbors will be removed by the thresholding, which we verify experimentally by finding almost no difference in PQ between $k=32$ and $64$. We thus keep $k=32$ for faster runtimes at a very little cost in terms of metrics.
5 Limitations and scope of this article
Since our method has a single fixed threshold per class, it is possible to craft failure cases by making two objects closer than the chosen threshold. It would thus be tempting to build intricate training-based solutions, for example by learning the neighboring condition [51, 52].
However, ALPINE proves that old techniques can reach state-of-the-art results, with no need for any object label, tuning, learning or extremely complex designs. ALPINE is a very strong baseline that future methods will have to beat to prove that their design actually benefits the task.
Furthermore, this object resolution limitation naturally reduces as the point density (or equivalently the LiDAR angular resolution) increases and intra-object average points distances reduces, leading to believe this limitation could mostly solve itself with new LiDAR technologies.
6 Conclusion
ALPINE is reaching state-of-the-art results in LiDAR panoptic segmentation with no instance labels, no training, and no heavy computation requiring a GPU. Furthermore, ALPINE can be taken off-the-shelf and applied as is on top of any semantic predictor.
Our oracle results indicate that improving semantic heads can yield significantly more PQ points than enhancing instance heads. Although this does not reflect the relative difficulty of such improvements, our study of ALPINE in replacement of prior instance predictions’ heads proves that instance extraction is largely saturated.
We conclude that panoptic labels are either not necessary, or not well utilized by current LiDAR panoptic segmentation methods. We encourage research to either focus more on semantics, or on how to better use instance annotations. We also release our code for integration into any semantic segmentation backbone for panoptic segmentation.
Supplementary Material
Appendix A Margin of improvement with oracle analysis
A.1 Discussion
Results in Sec. 4.5 motivate our conclusion that instance labels are mostly unnecessary or not correctly utilized for outdoor LiDAR panoptic segmentation.
On the other hand, we observe that the only methods competing with ALPINE on SemanticKITTI and nuScenes are end-to-end, query-based methods. This might hint that there is more to benefit from and end-to-end panoptic training than those two-staged oracles would suggest, for which we show in Tab. 7 that there is little gain left to expect with better instance predictions. Indeed end-to-end methods, and in particular query-based ones, benefit from the simultaneous instance and semantic segmentation tasks training. However, based on the tables of the main paper, the gain remains currently small, if positive, compared to our instance-annotation-free approach. This could hint at an existing margin for improvement for end-to-end methods.
Appendix B Metrics
In LiDAR point cloud segmentation, two primary metrics are used for evaluation: mean Intersection over Union (mIoU) and Panoptic Quality (PQ), each serving different segmentation goals.
Mean Intersection over Union (mIoU).
The mIoU is widely used in semantic segmentation to measure the overlap between predicted and ground truth masks for each class. For a given class $c$, the $\text{IoU}_{c}$ is defined as
$\text{IoU}_{c}=\dfrac{|O_{c}\cap G_{c}|}{|O_{c}\cup G_{c}|},$
where $O_{c}$ and $G_{c}$ represent predicted and ground truth masks for class $c$, respectively. The mIoU score is then averaged across all classes $C$:
$\text{mIoU}=\dfrac{1}{|C|}\sum_{c\in C}\text{IoU}_{c}.$
The mIoU does not differentiate between instances of the same class.
Panoptic Quality (PQ).
The PQ combines both semantic and instance segmentation and is computed as
$\text{PQ}_{c}=\underbrace{\frac{\sum_{(p,g)\in\text{TP}_{c}}\text{IoU}(p,g)}{|\text{TP}_{c}|}}_{\text{Segmentation Quality (SQ)}}\underbrace{\frac{|\text{TP}_{c}|}{|\text{TP}_{c}|\!+\!\tfrac{1}{2}|\text{FP}_{c}|\!+\!\tfrac{1}{2}|\text{FN}_{c}|}}_{\text{Recognition Quality (RQ)}},$
where $\text{TP}_{c}$, $\text{FP}_{c}$ and $\text{FN}_{c}$ are respectively the sets of true positives, false positives and false negatives computed after matching the predicted and ground truth instances in class $c$. The PQ score is then averaged across all classes:
$\text{PQ}=\frac{1}{|C|}\sum_{c\in C}\text{PQ}_{c}.$
[…]
$\text{PQ}^{\dagger}=\dfrac{1}{|\text{things}|\!+\!|\text{stuff}|}\left(\sum_{c\in\text{things}}\text{PQ}_{c}+\!\!\!\sum_{c\in\text{stuff}}\text{IoU}_{c}\right)\!.$
Appendix C Box fitting algorithm
Figure 6: Visual description of the box fitting algorithm
1 function fit_box(P)
     input: Points P
     output: Best fitting bounding box
[…]
3     $\texttt{B},\texttt{S}\leftarrow[],[]$
4     for $e\in E$ do
[…]
11          b($\texttt{b}\leftarrow(\texttt{b}^{\texttt{x}}_{\rm min},\texttt{b}^{\texttt{y}}_{\rm min},\texttt{b}^{\texttt{x}}_{\rm max},\texttt{b}^{\texttt{y}}_{\rm max})$
12          $\texttt{S}\texttt{.append}((\texttt{b}[2]-\texttt{b}[0])*(\texttt{b}[3]-\texttt{b}[1]))$
13          B.append(rotate(b, $\texttt{B}\texttt{.append}(\texttt{rotate(b, $-\boldsymbol{\theta}$)})$
14
[…]
16     return B[i]
Algorithm 2 Box fitting algorithm. This algorithm finds the best fitting box $B$ given a set of $2D$ points $P$. convex_hull returns the vertices and edges of the convex hull of P, angle gives the angle of an edge with respect to the $x$ axis, and rotate applies a rotation of a given angle.
The box fitting algorithm was borrowed from [78] which was inspired from [59]. It is described in Alg. 2 and Fig. 6. It works as follows. First, we compute the convex hull of the set of points on which we need to fit a box. Second, for each edge of the convex hull, we compute its angle with respect to the x-axis, we rotate the point cloud to align this edge with the x-axis, and then fit the axis-aligned bounding box of minimum area that cover the rotated point cloud. Finally, among all bounding boxes computed in the previous step (one for each edge of the convex hull), we keep the bounding box with the smallest area.
Appendix D Choice of margins in the box splitting
no split. 0%10%20%30%40%60%80%70707070727272727474747476767676787878788080808076.776.776.776.778.278.278.278.278.278.278.278.278.778.778.778.778.978.978.978.978.878.878.878.878.778.778.778.778.578.578.578.5Splitting marginPQ
Figure 7: Influence of the box splitting margin evaluated on SemanticKITTI’s validation set
[…]
First, we notice that the absence of margin hinders performance because of over-segmentation: no cluster bigger than the average box size can exist. Second, we remark that the results are stable for wide range of margins: only $0.2$ points of absolute variation in PQ for a margin between $20\%$ and $60\%$. Therefore, the margin does not need to be heavily finetuned.
The box splitting could reach a very high number of recursions, e.g., in the pathological case where points are to be removed one by one, or a high number of iterations in the dichotomy if the border between two objects is very sensitive. For this reason, we actually put a limit in how small $dt$ can be in Alg. 1. This limit is set to $31e-$ 3. This accelerates the algorithm by avoiding too long computations for corner cases. We verified in Tab. 10 that this choice does not impact the performance.
[…]
0.1 78.1 80.5 86.0
0.01 78.8 81.2 86.8
0.001 78.9 81.3 87.0
0.0001 78.9 81.3 87.0
0.00001 78.9 81.3 87.0
Table 10: Influence of the box dichotomy precision limit evaluated on nuScenes’ validation set.
Appendix E Explainability and limitations
ALPINE’s clustering is fully deterministic, and explainable in simple terms: in the absence of box splitting, two points $A$ and $B$ will belong to the same instance if they are predicted to be of the same semantics $c$, and there exist a path from $A$ to $B$ going only through points predicted to be of class $c$ with distance of at most $t_{c}$ between each edges. As such, a typical failure case exists when two instances of the same object are closer to each other than the threshold. The box splitting alleviates this issue but failures can still happen, e.g., if the distance between points within each instance is greater than the distance between the two instances. This issue vanishes as the angular resolution of the LiDAR is improved, which is experimentally verified by the higher PQ score achieved by a semantic oracle on the dense SemanticKITTI than on the sparser nuScenes.
Those weaknesses are a consequences of the simplicity and explainability of the method, and especially in the choice to use euclidean distance as the criterion of discrimination. ALPINE is a proof-of-concept of what can be achieved by a training-free method, and reaches state-of-the-art results, its focus is on being simple and lightweight, not on being all-encompassing.
Appendix F Per-class results
Method
barrier
bicycle
bus
car
con. veh.
motorcycle
pedestrian
tra. cone
PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU
ALPINE w. PTv3 61.9 84.4 77.3 54.0 83.0 96.1 93.3 95.2 64.4 49.8 89.9 89.7 93.8 86.6 92.4 75.0
ALPINE w. WI-768 61.0 84.3 77.5 54.8 81.0 94.5 93.2 95.4 61.1 49.4 90.8 90.1 93.6 86.2 92.6 76.2
Sem. Oracle 79.5 100 95.5 100 95.9 100 97.2 100 94.9 100 97.5 100 98.1 100 99.2 100
Method
trailer
truck
driveable
other flat
sidewalk
terrain
manmade
vegetation
PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU
ALPINE w. PTv3 66.2 78.6 72.8 86.6 57.6 76.2 87.7 89.8 97.0 97.2 63.6 75.4 90.0 91.9 73.1 77.0
ALPINE w. WI-768 64.6 79.4 75.2 87.0 57.3 76.6 85.9 88.8 96.9 97.1 58.8 75.0 89.0 91.1 72.3 76.4
Sem. Oracle 89.1 100 95.0 100 100 100 100 100 100 100 100 100 100 100 100 100
Table 11: Per-class panoptic segmentation results on nuScenes’ validation set.
Method
car
bicycle
motorcycle
truck
other-vehicle
person
bicyclist
motorcyclist
road
parking
PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU
ALPINE w. WI-256 93.0 96.9 68.4 61.9 79.4 84.7 72.2 92.2 61.3 66.4 87.4 83.0 91.7 93.2 00.0 00.7 95.7 95.7 35.6 51.1
ALPINE w. MinkUNet 93.1 98.0 69.5 64.8 81.1 87.0 79.0 93.7 73.6 84.9 87.7 83.3 93.9 94.3 12.6 23.4 94.3 94.7 42.9 53.7
Oracle 97.4 100 94.3 100 95.0 100 97.3 100 97.5 100 98.6 100 99.5 100 99.9 100 100 100 100 100
Method
sidewalk
other-ground
building
fence
vegetation
trunk
terrain
pole
traffic-sign
PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU
ALPINE w. WI-256 81.1 84.4 00.7 06.3 89.7 92.6 29.3 70.0 87.0 88.3 61.8 74.7 59.5 73.7 63.8 67.3 60.2 52.6
ALPINE w. MinkUNet 80.4 83.0 00.0 00.1 89.9 92.3 26.7 68.4 88.1 88.8 55.5 70.1 61.3 75.3 61.8 66.3 59.0 49.9
Oracle 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100
Table 12: Per-class panoptic segmentation results on SemanticKITTI’ validation set.
For completeness, we present in Tabs. 11 and 12 the per-class results of our method on both nuScenes and SemanticKITTI.
Appendix G More visualisations
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
D&M [84] w. MinkUNet ALPINE w. MinkUNet Ground Truth
Figure 8: Examples of panoptic predictions on SemanticKITTI. We present the results obtained with D&M [84] (left) and ALPINE (middle). Both methods use the same MinkUNet to obtain pointwise semantic predictions. The Ground Truth masks are presented on the rightmost panels. On the last row, only points predicted as “bicycle” are colored for visual clarity.
More visualization of our method can be seen on Fig. 8. We can see that D&M has some boundaries issues, which ALPINE does not have. Furthermore, on those examples, the remaining false predictions are mostly semantic prediction errors from the MinkUNet, rather than errors in our clustering.
[…]
labels
LPSAD (impl. from [51]) [39] 37.4 44.2 47.8 66.9 25.3 32.4 65.2 46.2 58.9 68.2 49.4
PanopticTrackNet [22] 40.0 - 48.3 73.0 29.9 33.6 76.8 47.4 59.1 70.3 53.8
PointGroup [23] 46.1 54.0 56.6 74.6 47.7 55.9 73.8 45.0 57.1 75.1 55.7
TORNADO-Net (fusion) [17] 50.6 55.9 62.1 74.9 48.1 57.5 72.5 52.4 65.4 76.7 59.2
Panoster [16] 55.6 - 66.8 79.9 56.6 65.8 - - - - 61.1
LCPS (lidar) [81] 55.7 65.2 65.8 74.0 - - - - - - 61.1
Cylinder3D & D&M [84] 57.2 - - - - - - - - - -
Location-Guided [69] 59.0 63.1 69.4 78.7 65.3 73.5 88.5 53.9 66.4 71.6 61.4
Panoptic-PolarNet [86] 59.1 64.1 70.2 78.3 65.7 74.7 87.4 54.3 66.9 71.6 64.5
EfficientLPS [58] 59.2 65.1 69.8 75.0 58.0 68.2 78.0 60.9 71.0 72.8 64.9
MaskPLS [35] 59.8 - 69.0 76.3 - - - - - - 61.9
DS-Net (SPVCNN) [21] 61.4 65.2 72.7 79.0 65.2 72.3 79.3 57.9 71.1 79.3 69.6
Panoptic-PHNet [30] 61.7 - - - 69.3 - - - - - 65.7
PANet [36] 61.7 66.6 71.8 79.3 - - - - - - 68.1
CenterLPS [37] 62.1 67.0 72.0 80.7 68.4 75.2 91.0 57.5 69.7 73.2 68.1
P3Former [70] 62.6 66.2 72.4 76.2 69.4 - - 57.7 - - -
CFNet [31] 62.7 67.5 - - 70.0 - - 57.3 - - 67.4
LPST [2] 63.1 70.8 73.1 79.2 68.7 75.7 86.7 58.9 71.2 73.7 69.7
GP-S3Net [51] 63.3 71.5 75.9 81.4 70.2 80.1 86.2 58.3 72.9 77.9 73.0
DQFormer [76] 63.5 67.2 73.1 81.7 - - - - - - -
ALPINE w. MinkUNet 64.2 68.9 74.1 84.4 72.6 79.3 90.6 58.1 70.4 79.8 70.7
PUPS (w/o ens.) [60] 64.4 68.6 74.1 81.5 73.0 79.3 92.6 58.1 70.4 73.5 -
IEQLPS [19] 64.7 68.1 74.7 81.3 73.5 79.1 93.0 58.3 71.4 72.8 -
Methods with Test-Time Augmentations and/or model ensembling
D&M w. MinkUnet [84] (✓) 61.8 66.2 72.8 79.6 64.5 73.7 87.2 59.9 72.1 74.2 71.4
ALPINE w. PTv3 (✓) 62.4 66.2 72.0 76.7 66.1 72.1 80.0 59.8 71.9 74.3 67.3
ALPINE w. WI-256 ** (✓) 64.2 69.0 74.1 79.7 69.3 75.9 79.8 60.4 72.7 79.7 70.3
ALPINE w. MinkUNet (✓) 65.9 70.2 75.5 81.4 73.9 80.1 91.2 60.0 72.2 74.2 72.2
PUPS [60] 66.3 70.2 75.6 82.5 74.6 80.3 93.4 60.2 72.2 74.5 -
ALPINE w. MinkUNet & PTv3 (✓) (✓) 66.6 70.8 76.1 82.6 73.4 79.5 93.2 61.6 73.6 74.9 72.0
*: the publicly available model for WI-256 [45] used instance annotations in its data augmentation pipeline
Table 13: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
[…]
labels
LPSAD (impl. from [51]) [39] 50.4 57.7 62.4 79.4 43.2 53.2 80.2 57.5 71.7 78.5 62.5
PanopticTrackNet [22] 51.4 56.2 63.3 80.2 45.8 55.9 81.4 60.4 75.5 78.3 58.0
VIN [85] 51.7 57.4 61.8 82.6 45.7 53.7 83.6 61.8 75.4 80.9 73.7
TORNADO-Net (fusion) [17] 54.0 59.8 65.4 80.9 44.1 53.9 80.1 63.9 76.9 81.8 68.0
MaskPLS [35] 57.7 60.2 66.0 71.8 64.4 73.3 84.8 52.2 60.7 62.4 62.5
GP-S3Net [51] 61.0 67.5 72.0 84.1 56.0 65.2 85.3 66.0 78.7 82.9 75.8
EfficientLPS [58] 62.0 65.6 73.9 83.4 56.8 68.0 83.2 70.6 83.6 83.8 65.6
DS-Net (SPVCNN) [21] 64.7 67.6 76.1 83.5 58.6 64.2 82.8 74.7 86.5 85.5 76.3
PVCL [32] 64.9 67.8 77.9 81.6 59.2 72.5 79.7 67.6 79.1 77.3 73.9
SCAN [73] 65.1 68.9 75.3 85.7 60.6 70.2 85.7 72.5 83.8 85.7 77.4
Panoptic-PolarNet [86] 67.7 71.0 78.1 86.0 65.2 74.0 87.2 71.9 84.9 83.9 69.3
SMAC-Seg [27] 68.4 73.4 79.7 85.2 68.0 77.2 87.3 68.8 82.1 83.0 71.2
PANet [36] 69.2 72.9 80.7 85.0 69.5 79.3 86.7 68.7 82.9 82.1 72.6
CPSeg [28] 71.1 75.6 82.5 85.5 71.5 81.3 87.3 70.6 83.7 83.6 73.2
LCPS (lidar) [81] 72.9 77.6 82.0 88.4 72.8 80.5 90.1 73.0 84.5 85.5 75.1
PUPS [60] 74.7 77.3 83.3 89.4 75.4 81.9 91.8 73.6 85.6 85.3 -
Panoptic-PHNet [30] 74.7 77.7 84.2 88.2 74.0 82.5 89.0 75.9 86.9 86.8 79.7
CFNet [31] 75.1 78.0 84.6 88.8 74.8 82.9 89.8 76.6 87.3 87.1 79.3
P3Former [70] 75.9 78.9 84.7 89.7 76.9 83.3 92.0 75.4 87.1 86.0 -
CenterLPS [37] 76.4 79.2 86.2 88.0 77.5 87.1 88.4 74.6 84.9 87.3 77.1
ALPINE w. WI-768 76.9 79.9 85.7 89.3 77.9 85.3 90.9 75.3 86.3 86.7 80.3
IEQLPS [19] 77.1 79.1 86.5 88.2 79.5 86.6 91.7 73.0 86.4 83.9 -
LPST [2] 77.1 79.9 86.5 88.6 79.3 87.5 90.3 73.6 84.9 85.7 80.3
DQFormer [76] 77.7 79.5 86.8 89.2 77.8 86.7 89.5 77.5 87.0 88.6 -
Methods with Test-Time-Augmentations and/or model ensembling
ALPINE w. WI-768 (✓) 77.9 80.7 86.5 89.6 78.6 86.0 91.0 76.7 87.3 87.3 81.4
ALPINE w. PTv3 (✓) 78.9 81.3 87.0 90.4 79.3 86.2 91.7 78.2 88.3 88.1 81.5
ALPINE w. WI-768 & PTV3 (✓) (✓) 79.5 81.9 87.6 90.5 80.1 87.1 91.7 78.6 88.5 88.3 82.7
LidarMultiNet [77] 81.8 - 89.7 90.8 - - - - - - 83.6
Table 14: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
[…]
labels
LPSAD (impl. from [51]) [39] 22.5 32.7 34.0 53.5 18.7 25.7 70.5 24.0 37.1 47.1 35.5
TORNADO-Net (fusion) [17] 33.7 43.3 46.0 68.4 41.2 49.6 83.1 30.9 44.7 62.9 44.5
DS-Net [20] 35.6 45.9 49.2 68.6 27.4 33.8 76.8 38.7 55.0 65.5 54.5
GP-S3Net [51] 48.7 60.3 63.7 61.3 61.6 71.7 86.4 43.8 60.8 51.8 61.8
ALPINE w. PTv3 51.4 57.7 67.2 74.3 64.8 74.1 87.2 47.4 65.1 70.5 58.3
Table 15: Panoptic segmentation results on Sequence $02$ of SemanticPOSS as validation set.

### v2

@@
 In this work, we demonstrate that competitive panoptic segmentation can be achieved using only semantic labels, with instances predicted without any training or annotations.
-Our method achieves performance comparable to current state-of-the-art supervised methods on standard benchmarks including SemanticKITTI and nuScenes, and outperforms every publicly available method on SemanticKITTI as a drop-in instance head replacement, while running in real-time on a single-threaded CPU and requiring no instance labels.
-Our method is fully explainable, and requires no learning or parameter tuning.
+Our method outperforms state-of-the-art supervised methods on standard benchmarks including SemanticKITTI and nuScenes, and outperforms every publicly available method on SemanticKITTI as a drop-in instance head replacement, while running in real-time on a single-threaded CPU and requiring no instance labels.
+It is fully explainable, and requires no learning or parameter tuning.
+ALPINE combined with state-of-the-art semantic segmentation ranks first on the official panoptic segmentation leaderboard of SemanticKITTI.
 Code is available at https://github.com/valeoai/Alpine/
@@
 lbs.
-LPSAD (in [51]) [39] 37.4 44.2 47.8 66.9 49.4
-Panop.TrackNet [22] 40.0 - 48.3 73.0 53.8
-PointGroup [23] 46.1 54.0 56.6 74.6 55.7
-TORNADO-Net [17] 50.6 55.9 62.1 74.9 59.2
-Panoster [16] 55.6 - 66.8 79.9 61.1
 LCPS (lidar) [82] 55.7 65.2 65.8 74.0 61.1
 D&M (Cyl.3D) [85] 57.2 - - - -
-Location-Guided [69] 59.0 63.1 69.4 78.7 61.4
 Panop.-PolarNet [87] 59.1 64.1 70.2 78.3 64.5
-EfficientLPS [58] 59.2 65.1 69.8 75.0 64.9
 MaskPLS [36] 59.8 - 69.0 76.3 61.9
@@
 Table 1: Panoptic segmentation results on the validation set of SemanticKITTI.
-‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively.
-(✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
+‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling.
+(✓) denotes that TTA/ensemble was used on the semantic head only.
 The main metric is the PQ.
+A more detailed version of this table is available in the supplementary material.
 {NiceTabular}
@@
 lbs.
-LPSAD (in [51]) [39] 50.4 57.7 62.4 79.4 62.5
-Panop.TrackNet [22] 51.4 56.2 63.3 80.2 58.0
-VIN [85] 51.7 57.4 61.8 82.6 73.7
-TORNADO-Net [17] 54.0 59.8 65.4 80.9 68.0
-MaskPLS [35] 57.7 60.2 66.0 71.8 62.5
-GP-S3Net [51] 61.0 67.5 72.0 84.1 75.8
-EfficientLPS [58] 62.0 65.6 73.9 83.4 65.6
 DS-Net(SPVCNN) [21] 64.7 67.6 76.1 83.5 76.3
-PVCL [32] 64.9 67.8 77.9 81.6 73.9
-SCAN [73] 65.1 68.9 75.3 85.7 77.4
 Panop.-PolarNet [87] 67.7 71.0 78.1 86.0 69.3
-SMAC-Seg [27] 68.4 73.4 79.7 85.2 71.2
 PANet [37] 69.2 72.9 80.7 85.0 72.6
@@
 Table 2: Panoptic segmentation results on the validation set of nuScenes.
-‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively.
-(✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
+‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling.
+(✓) denotes that TTA/ensemble was used on the semantic head only.
 The main metric is the PQ.
+A more detailed version of this table is available in the supplementary material.
 {NiceTabular}
@@
 4.3 Comparison to the state-of-the-art
+[…]
+5/ P3Former 65.3 67.8 74.9 86.6 67.4 66.1
+4/ PUPS 65.7 70.3 75.8 85.7 68.1 68.5
+3/ P3Former & ALPINE 66.1 68.7 75.9 86.7 69.4 66.1
+2/ UniSeg 67.2 72.1 78.1 85.5 67.5 73.8
+1/ UniSeg & ALPINE 70.2 75.1 80.5 86.7 74.3 75.1
+Table 4: SemanticKITTI test set results (official leaderboad on June 12, 2025). We used P3Former’s and UniSeg’s semantic predictions and used ALPINE for instance segmentation.
 sem. Original head ALPINE FPS
@@
 PQ FPS PQ FPS PQ FPS PQ FPS
-65.5 14.4 Hz 56.7 3.2 Hz 55.1 4.5 61.8 0.5 Hz
+65.5 14.4 Hz 56.7 3.2 Hz 55.1 4.5 Hz 61.8 0.5 Hz
 […]
@@
 Comparison to other instance heads.
+We present results obtained with ALPINE on the test set of SemanticKITTI.
+First, we use the public model of P3Former for semantic prediction and extract object instances using ALPINE.
+The PQ which is at 65.3 with P3Former improves to 66.1 thanks to ALPINE, the third best entry on the leaderboard.
+Second, we repeat the same procedure with the semantic labels of UniSeg [35] (labels kindly provided by the authors).
+UniSeg with ALPINE’s instance labels ranks first on the leaderboard of SemanticKITTI.
+These results highlight the high quality of the instances extracted by our method.
+We note nevertheless that these results are not strictly obtained without instance labels as both UniSeg and P3Former did use them during training.
+We also remark that while the semantic mIoU of P3Former and P3Former&ALPINE are identical, UniSeg&ALPINE’s semantic mIoU is above UniSeg’s: 75.1 vs 73.8.
+We did not modify UniSeg’s semantic labels (75.1 is in line with the score of UniSeg on the semantic segmentation leaderboard).
+We suspect a misformatting in UniSeg’s panoptic labels but which we cannot correct ourselves.
+We conduct a more complete study on the validation set of SemanticKITTI.
 Tab.
@@
 This approximate measurement is comparable (while not directly equivalent) to the ALPINE FPS reported in Tab.
-5.
+6.
 Comparison to other clustering methods.
@@
 As can be noted in Tab.
-6, ALPINE significantly outperforms those methods.
+7, ALPINE significantly outperforms those methods.
 It can still be noted that both unsupervised method work quite impressively, almost reaching the level of performance of D&M while not using any semantic information.
@@
 Results, in Tab.
-7, shows how much scores are expected to improve with both oracles.
+8, shows how much scores are expected to improve with both oracles.
 When using the “semantic oracle”, the PQ on nuScenes raises to 96.2, (and 99.0 on SemanticKITTI), while using PTv3 predictions and the instance oracle only raises the PQ by +4.3 to 83.2, showing that instance extraction is much more saturated than semantic segmentation.
@@
 PTv3 58.3 51.1 57.4 63.6 51.4 57.7 64.8
-Table 9: Comparison between statistics-based and web-based parameters settings. More run details can be found in Secs. 4 and 4.
+Table 10: Comparison between statistics-based and web-based parameters settings. Run details can be found in Secs. 4 and 4.
 4.6 Ablation and sensitivity study
@@
 We present our ablation study in Tab.
-8, conducted on the validation set of nuScenes, using PTv3 as semantic backbone.
+9, conducted on the validation set of nuScenes, using PTv3 as semantic backbone.
 We see that constructing the graph in BEV or post-processing the result with our box splitting strategy improve the results: respectively $+0.6$ point and $+0.8$ point in PQ over the baseline.
@@
 We also release our code for integration into any semantic segmentation backbone for panoptic segmentation.
+Acknowledgements.
+We wish to thank the authors of P3Former [71] for making their code and model used on the test set of SemanticKITTI available.
+We also thank the authors of UniSeg [35], particularly Lingdong Kong, for providing us with their test set semantic predictions.
+This work was supported in part by the French Agence Nationale de la Recherche (ANR) grant MultiTrans (ANR21-CE23-0032).
 Supplementary Material
@@
 Table 13: Per-class panoptic segmentation results on SemanticKITTI’ validation set.
-For completeness, we present in Tabs. 11 and 12 the per-class results of our method on both nuScenes and SemanticKITTI.
-Appendix G More visualisations
+For completeness, we present in Tabs. 12 and 13 the per-class results of our method on both nuScenes and SemanticKITTI.
+Appendix G Performance of ALPINE and distance to the sensor
+0-15m 15-30m 30m+
+[…]
+ALPINE w/ Ens. 68.5 71.0 73.0 68.6 70.7 68.4 65.8 65.9 54.4
+P3Former 63.6 65.8 67.8 64.8 66.9 63.4 59.6 60.1 47.2
+P3Former & ALPINE 64.0 66.2 67.8 64.8 66.9 63.4 59.4 59.8 47.2
+Table 14: Performance w.r.t sensor distance on SemanticKITTI’s validation set.
+We report in Tab.
+14 a breakdown of ALPINE’s performance with respect to the distance to the sensor.
+ALPINE improves over P3Former at short range, but slightly deteriorates at long range.
+This suggests that ALPINE is particularly efficient on dense point clouds (i.e.
+close to the sensor) but gets lesser performance on sparser point clouds.
+Appendix H More visualisations
 Refer to caption Refer to caption Refer to caption
@@
 Table 15: Panoptic segmentation results on the validation set of SemanticKITTI.
-‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively.
-(✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
+‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling.
+(✓) denotes that TTA/ensemble was used on the semantic head only.
+The main metric is the PQ.
 […]
@@
 Table 16: Panoptic segmentation results on the validation set of nuScenes.
-‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively.
-(✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
+‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling.
+(✓) denotes that TTA/ensemble was used on the semantic head only.
+The main metric is the PQ.
 […]